OpenStudy (raylenesolis):

Describe how to transform the quantity of the fifth root of x to the seventh power, to the third power into an expression with a rational exponent. Make sure you respond with complete sentences.

jhonyy9 (jhonyy9):

so fifth root of x mean x^(1/5) yes ?

jhonyy9 (jhonyy9):

and for x on 7 power will be (x^7)^(1/5) = x^(7/5)

jhonyy9 (jhonyy9):

and how is about this 3rd power again ?

OpenStudy (raylenesolis):

\[(\sqrt[5]{x ^{7}})^{3}\] it looks like this

TheSmartOne (thesmartone):

Hint: \(\Large \sqrt[c]{a^b} = a^{\frac{b}{c}}\) \(\Large (x^y)^z = x^{yz}\)